# Introducing Perturb-ability Score (PS) to Enhance Robustness Against Evasion Adversarial Attacks on ML-NIDS Author’s draft for soliciting feedback: 5 Nov. 2024

### v2

@@ 8 Conclusion
This paper investigates the perturb-ability of features within NIDS by introducing the Perturb-ability Score (PS) metric. A high PS indicates that a feature is easily manipulated in the problem-space while maintaining the malicious and network functionality of the modified flow, while a low PS suggests that altering that feature may either be infeasible or could render the network flow or malicious objective invalid. Through our research, we have leveraged the PS as a potential defense mechanism against practical problem-space evasion adversarial attacks by selectively identifying and utilizing features with low PS. This novel strategy introduces a fresh approach, utilizing NIDS domain constraints to mitigate the impact of adversarial attacks on ML-based network intrusion detection systems.
The results from our experiments demonstrate that discarding features with high PS does not compromise model performance. Instead, it enhances the robustness of NIDS against adversarial threats, providing reliable detection capability. The findings suggest that by focusing on inherently robust features, we can effectively fortify NIDS against problem-space evasion adversarial attacks while maintaining high detection accuracy.
Overall, this work presents a promising avenue for enhancing the security of ML-based NIDS. Future research will focus on further refining the PS metric and exploring its applicability across different datasets and attack scenarios, ultimately contributing to the development of more resilient network security systems.
@@ Appendix A Definition of Features of UNSW-NB15 and improved CSE-CIC-IDS2018 Datasets
TABLE VI: The Features Description of the UNSW-NB15 Dataset
@@
TABLE VII: The Features Description of the CSE-CIC-IDS2018 Dataset (Part A)
@@
TABLE VIII: The Features Description of the CSE-CIC-IDS2018 Dataset (Part B)

### v3

@@ 8 Conclusion
This paper investigates the perturb-ability of features within NIDS by introducing the Perturb-ability Score (PS) metric. A high PS indicates that a feature is easily perturb-able in the problem-space while maintaining NIDS domain constrains, while a low PS suggests that altering that feature may either be infeasible or could render the network flow or malicious objective invalid. Through our research, we have leveraged the PS as a potential defense mechanism against practical problem-space evasion adversarial attacks by selectively identifying and utilizing features with low PS. This novel strategy introduces a fresh approach, utilizing NIDS domain problem-space constraints to mitigate the impact of adversarial attacks on Flow-based ML-NIDS. The results from our experiments demonstrate that discarding features with high PS does not compromise model performance. Instead, it enhances the robustness of NIDS against adversarial threats, providing reliable detection capability. The findings suggest that by focusing on inherently robust features, we can effectively fortify NIDS against problem-space evasion adversarial attacks while maintaining high detection accuracy. Overall, this work presents a promising avenue for enhancing the security of ML-based NIDS. Future research will focus on further refining the PS metric and exploring its applicability across different datasets and attack scenarios, ultimately contributing to the development of more resilient network security systems.
@@ Appendix A Definition of Features of UNSW-NB15 and improved CSE-CIC-IDS2018 Datasets
Table 6: The Features Description of the UNSW-NB15 Dataset
@@
Table 7: The Features Description of the CSE-CIC-IDS2018 Dataset (Part A)
@@
Table 8: The Features Description of the CSE-CIC-IDS2018 Dataset (Part B)